# Motivic Gauss and Jacobi sums

### v2

@@ 1. Introduction
$g_{d}=-\sum_{m\in\kappa^{*}}(m,m^{\frac{q-1}{d}})\in\mathbb{Z}[\kappa\times\mu_{d}]$
induces the Frobenius endomorphism on $h(A_{d})^{(\psi,\chi)}$ if $\psi$ and $\chi$ are non-trivial. Since the Frobenius endomorphism commutes with any morphisms by [9, p. 80] (cf. [20, Proposition 2]), (1.1) and (1.2) can be deduced from Coleman’s result and Theorems 1.1, 1.2.
We conclude this introduction by a discussion on the relations among Weil numbers and motives. Recall that $\alpha\in\mathbb{C}$ is called a $q$-Weil number of weight $w\in\mathbb{Z}$ if there exists $m\in\mathbb{Z}$ such that $q^{m}\alpha$ is an algebraic integer and $|\sigma(\alpha)|=q^{w/2}$ for all $\sigma\colon\mathbb{Q}(\alpha)\to\mathbb{C}$. Let $W_{q}(\Lambda)$ be the subgroup of $\Lambda^{*}$ consisting of all $q$-Weil numbers (of arbitrary weight) belonging to a given subfield $\Lambda$ of $\mathbb{C}$. It is conjectured by Beilinson [2, 1.0] that the rational equivalence and numerical equivalence should agree over a finite field $\kappa$ (with coefficients in $\Lambda$). If we assume this as well as the Tate conjecture, it follows from [12, Proposition 2.21] that all simple objects of $\mathbf{Chow}(\kappa,\overline{\mathbb{Q}})$ should be invertible, and their isomorphism classes would form a group (with respect to the tensor product) isomorphic to $W_{q}(\overline{\mathbb{Q}})$. In particular, the group $\mathbf{Pic}(\mathbf{Chow}(\kappa,\Lambda))$ of all isomorphism classes of invertible objects of $\mathbf{Chow}(\kappa,\Lambda)$ should be isomorphic to a subgroup of $W_{q}(\Lambda)$ by Lemma 3.3 below. Therefore, the relations among $q$-Weil numbers (such as (1.1) and (1.2)) should come from relations among motives, as demonstrated by Theorem 1.1 (which holds unconditionally and over any $\kappa$).
@@ 3.1. Chow motives
Let $f\colon X\to Y$ be a generically finite morphism of degree $d$, where $X,Y\in\mathbf{SmProj}(\kappa)$ are both equi-dimensional of the same dimension $m$. We define a group homomorphism
@@
    Let $\kappa$ be a finite field with $q$ elements and $\mathrm{Fr}_{X}$ be the $q$th power Frobenius endomorphism of $X$. Then $f_{\#}(\mathrm{Fr}_{X})=\mathrm{Fr}_{Y}$ (cf. [9, p. 80], [20, Proposition 2]).
@@
Let Λsuperscriptnormal-$\Lambda^{\prime}$ be a field extension of $\Lambda$. Then the scalar extension functor $\mathbf{Chow}(\kappa,\Lambda)\to\mathbf{Chow}(\kappa,\Lambda^{\prime})$ is conservative.
@@ 3.2. Voevodsky motives
[…]
@@ 4.1. Artin-Schreier motives
Here, $e^{(\psi,\chi)}$ means the algebraic correspondence induced by the group-ring element defined in (2.8). Define a projector $e_{\mathrm{prim}}\in\mathbb{Q}[\kappa\times\mu_{d}]$ by
@@
Put $h(A_{d})_{\mathrm{prim}}=(A_{d},e_{\mathrm{prim}})=\bigoplus_{\psi\neq 1,\chi\neq 1}h(A_{d})^{(\psi,\chi)}$.
@@
Suppose that the intersection number $n:=(e,e^{\prime})$ is invertible in $\Lambda$, and define
@@
[…]
@@ 4.2. Fermat motives
and define a projector $e_{\mathrm{prim}}\in\mathbb{Q}[\mu_{d}^{n}]$ by
@@
Hence the first assertion follows. Since $h(F_{d}^{(1)})\simeq\Lambda[\mu_{d}]$, we have $h(F_{d}^{(1)})^{\chi}\simeq\Lambda$, and the second assertion follows. ∎
@@
[…]
@@
[…]
@@ 5.2. Proof of the case $n=2$
By Propositions 4.1 and 4.6, we can suppose $d=q-1$, so that $\mu_{d}=\kappa^{*}$. Here we need $\mathbf{DM}_{\mathrm{gm}}(\kappa,\Lambda)$ from subsection 3.2 to treat open varieties. We just write $A=A_{q-1}$, $F=F_{q-1}^{(2)}$. Let $A^{\circ}\subset A$ (resp. $F^{\circ}\subset F$) be the affine open subscheme defined in (4.1) (resp. by $u_{0}\neq 0$). Write $A^{\circ}F=A^{\circ}\times F$, $A^{\circ}F^{\circ}=A^{\circ}\times F^{\circ}$. Define a correspondence $\Gamma\subset(A^{\circ})^{2}\times A^{\circ}F$ by
@@ 7.2. Multiplication formula
One recovers (7.5) from the corollary and the invertibility of the motives, noting
@@
If $\kappa=\mathbb{C}$, the complex period of an invertible motive in $\mathbf{Chow}(\mathbb{C},\mathbb{Q}(\zeta_{d}))$ is an element of *$\mathbb{C}^{*}/\mathbb{Q}(\zeta_{d})^{*}$, defined by the de Rham-Betti comparison isomorphism. Since the periods of Fermat motives are special values of the beta function, the isomorphism (7.6) implies (7.3) for any $s\in d^{-1}\mathbb{Z}$, up to $\mathbb{Q}(\zeta_{d})^{*}$.
@@ 8. Weil numbers
where $a_{0}:=-\sum_{i=1}^{n}a_{i}$. If $\mathbf{a}\in A_{d}^{(n)}$, then $j_{d}(\mathbf{a})$ is a $q$-Weil number of weight $n-1$ by (2.6). Let $W=W_{q}(F)\subset F^{*}$ be the subgroup of $q$-Weil numbers and $J\subset W$ be the subgroup generated by $\{j_{d}(\mathbf{a})\mid\mathbf{a}\in A_{d}^{(n)},n\geq 2\}$. Since $\sigma_{h}j_{d}(\mathbf{a})=j_{d}(h\mathbf{a})$, $J$ is $G$-stable as well as $W$. In fact, $J$ is generated by $\{j_{d}(\mathbf{a})\mid\mathbf{a}\in A_{d}^{(2)}\}\cup\{(-1)^{\frac{q-1}{d}}\}$ by (6.3).
@@
which associates to an invertible motive its Frobenius eigenvalue. As a result of Theorem 1.2 and Proposition 8.2, we obtain the following.
@@
Assume the conjectures of Beilinson and Tate. Then $\mathbf{Pic}(\mathbf{Chow}(\kappa,F))_{\mathbb{Q}}$ is isomorphic to $W_{\mathbb{Q}}$ and is generated as a $\mathbb{Q}$-vector space by the classes of Fermat motives $h(F_{d}^{(2)})^{{\boldsymbol{\chi}}}$ (${\boldsymbol{\chi}}\in\mathfrak{X}_{d}^{(2)}$).

### v3

@@ 1. Introduction
(1.3) $g_{d}=-\sum_{m\in\kappa^{*}}(m,m^{\frac{q-1}{d}})\in\mathbb{Z}[\kappa\times\mu_{d}]$
induces the Frobenius endomorphism on $h(A_{d})^{(\psi,\chi)}$ if $\psi$ and $\chi$ are non-trivial. Since the Frobenius endomorphism commutes with any morphisms (see (3.4) below), (1.1) and (1.2) can be deduced from Coleman’s result and Theorems 1.1, 1.2. See Remark 7.4 for details.
We conclude this introduction by a discussion on the relations among Weil numbers and motives. Recall that $\alpha\in\mathbb{C}$ is called a $q$-Weil number of weight $w\in\mathbb{Z}$ if there exists $m\in\mathbb{Z}$ such that $q^{m}\alpha$ is an algebraic integer and $|\sigma(\alpha)|=q^{w/2}$ for all $\sigma\colon\mathbb{Q}(\alpha)\to\mathbb{C}$. Let $W_{q}(\Lambda)$ be the subgroup of $\Lambda^{*}$ consisting of all $q$-Weil numbers (of arbitrary weight) belonging to a given subfield $\Lambda$ of $\mathbb{C}$. It is conjectured by Beilinson [2, 1.0] that the rational equivalence and numerical equivalence should agree over a finite field $\kappa$ (with coefficients in $\Lambda$). If we assume this as well as the Tate conjecture, it follows from [12, Proposition 2.21] that all simple objects of $\mathbf{Chow}(\kappa,\overline{\mathbb{Q}})$ should be invertible, and their isomorphism classes would form a group (with respect to the tensor product) isomorphic to $W_{q}(\overline{\mathbb{Q}})$. In particular, the group $\mathbf{Pic}(\mathbf{Chow}(\kappa,\Lambda))$ of all isomorphism classes of invertible objects of $\mathbf{Chow}(\kappa,\Lambda)$ should be isomorphic to a subgroup of $W_{q}(\Lambda)$ by Lemma 3.3 below. Therefore, the multiplicative relations among $q$-Weil numbers (such as (1.1) and (1.2)) should come from relations among motives, as demonstrated by Theorem 1.1.
@@
Acknowledgements
The authors thank sincerely the referee for his/her careful reading and helpful comments.
@@ 3.1. Chow motives
Let $f\colon X\to Y$ be a generically finite morphism of degree $d$, where $X,Y\in\mathbf{SmProj}(\kappa)$ are both irreducible and of the same dimension $m$. We define a group homomorphism
@@
    Let $\kappa$ be a finite field with $q$ elements and $\mathrm{Fr}_{X}$ be the $q$th power Frobenius endomorphism of $X$. Then $f_{\#}(\mathrm{Fr}_{X})=\mathrm{Fr}_{Y}$ holds, because we have
    (3.4) $f\circ\mathrm{Fr}_{X}=\mathrm{Fr}_{Y}\circ f$
    for any morphism $f\colon X\to Y$ in $\mathbf{Chow}(\kappa,\Lambda)$ by [9, p. 80] (see also [20, Proposition 2]).
@@
Let $\Lambda^{\prime}$ be a field extension of $\Lambda$. Then the scalar extension functor $\mathbf{Chow}(\kappa,\Lambda)\to\mathbf{Chow}(\kappa,\Lambda^{\prime})$ is conservative.
@@ 3.2. Voevodsky motives
$f_{*}\circ f^{*}=(\deg f)\cdot\mathrm{id}_{V}\quad\text{in}\ \mathbf{Cor}(V,V),\quad f^{*}\circ f_{*}=\sum_{g\in\mathrm{Gal}(U/V)}g_{*}\quad\text{in}\ \mathbf{Cor}(U,U).$
@@ 4.1. Artin-Schreier motives
Here, $e^{(\psi,\chi)}$ means the algebraic correspondence induced by the group-ring element defined in (2.8). Define a projector $e_{\mathrm{prim}}\in\mathbb{Q}[\kappa\times\mu_{d}]$ (with coefficients in $\mathbb{Q}$) by
@@
Note that, for any projectors $e,f\in\mathbb{Q}[G]$ where $G$ is an abelian group, $1-e$ and $ef$ are also projectors. Put $h(A_{d})_{\mathrm{prim}}=(A_{d},e_{\mathrm{prim}})=\bigoplus_{\psi\neq 1,\chi\neq 1}h(A_{d})^{(\psi,\chi)}$.
@@
Suppose that the intersection number $n:=(e,e^{\prime})$ is not zero, and define
@@
    $\underline{\alpha}\circ e\colon\Lambda(1)\to M\quad\text{and}\quad\frac{1}{n}e^{\prime}\circ\overline{\alpha}\colon M\to\Lambda(1)$
@@ 4.2. Fermat motives
and define a projector $e_{\mathrm{prim}}\in\mathbb{Q}[\mu_{d}^{n}]$ (with coefficients in $\mathbb{Q}$) by
@@
Hence the first assertion follows. Since $h(F_{d}^{(1)})\simeq\Lambda[\mu_{d}]$, the Artin motive of the regular representation of $\mu_{d}$, we have $h(F_{d}^{(1)})^{\chi}\simeq\Lambda$, and the second assertion follows. ∎
@@
    $h(Z_{1})^{\boldsymbol{\chi}}\simeq\begin{cases}\Lambda\langle-1\rangle^{\chi_{n}}&\text{if $\chi_{1}=\cdots=\chi_{n-1}=\chi_{n}\chi_{n+1}=1$},\\ 0&\text{otherwise}.\end{cases}$
    $h(Z_{2})^{\boldsymbol{\chi}}\simeq\begin{cases}h(F_{d}^{(n-1)})^{(\chi_{1},\dots,\chi_{n-1})}&\text{if $\chi_{n}=\chi_{n+1}=1$},\\ 0&\text{otherwise}.\end{cases}$
@@
    $h(F_{d}^{(n,2)}/H)^{{\boldsymbol{\chi}}^{\prime}}\simeq\begin{cases}h(F_{d}^{(n,2)})^{{\boldsymbol{\chi}}^{\prime}}&\text{if $\chi_{n}=\chi_{1}^{\prime}\chi_{2}^{\prime}$},\\ 0&\text{otherwise}.\end{cases}$
    $h(Z)^{{\boldsymbol{\chi}}^{\prime}}\simeq\begin{cases}h(F_{d}^{(n-1)})^{(\chi_{1},\dots,\chi_{n-1})}\otimes\Lambda\langle-1\rangle^{\chi_{1}^{\prime}}&\text{if $\chi_{n}=\chi_{1}^{\prime}\chi_{2}^{\prime}=1$},\\ 0&\text{otherwise}.\end{cases}$
@@ 5.2. Proof of the case $n=2$
By Propositions 4.1 and 4.6, we can suppose $d=q-1$, so that $\mu_{d}=\kappa^{*}$. Here we need $\mathbf{DM}_{\mathrm{gm}}(\kappa,\Lambda)$ from subsection 3.2 to treat open varieties. We just write $A=A_{q-1}$, $F=F_{q-1}^{(2)}$. Let $A^{\circ}\subset A$ (resp. $F^{\circ}\subset F$) be the affine open subscheme defined in (4.1) (resp. by $u_{0}\neq 0$). Write $A^{\circ}F=A^{\circ}\times F$, $A^{\circ}F^{\circ}=A^{\circ}\times F^{\circ}$. Define a closed subscheme $\Gamma\subset(A^{\circ})^{2}\times A^{\circ}F$ by
@@ 7.2. Multiplication formula
Remark 7.4.
One can deduce (1.1) and (1.2) from Coleman’s theorem (1.3) and Theorems 1.1, 1.2 as follows. Since both sides of Theorem 1.1 (i) are invertible, their endomorphism rings are canonically isomorphic to $\Lambda$. The Frobenius endomorphisms on both sides, regarded as elements of $\Lambda$, yield the same element because the Frobenius endomorphism commutes with any morphism (see (3.4)). We conclude (1.1) by observing that the Frobenius endomorphisms agree with the left and right hand sides of (1.1) by (1.3) and Theorem 1.2, respectively.
Similarly, one recovers (1.2) (which is (7.5)) from Corollary 7.3 and the invertibility of the motives, noting
@@
If $\kappa=\mathbb{C}$, the complex period of an invertible motive in $\mathbf{Chow}(\mathbb{C},\mathbb{Q}(\zeta_{d}))$ is an element of $\mathbb{C}^{*}/\mathbb{Q}(\zeta_{d})^{*}$, defined by the de Rham-Betti comparison isomorphism. Since the periods of Fermat motives are special values of the beta function, the isomorphism (7.6) implies (7.3) for any $s\in d^{-1}\mathbb{Z}$, up to $\mathbb{Q}(\zeta_{d})^{*}$.
@@ 8. Weil numbers
where $a_{0}:=-\sum_{i=1}^{n}a_{i}$. If $\mathbf{a}\in A_{d}^{(n)}$, then $j_{d}(\mathbf{a})$ is a $q$-Weil number of weight $n-1$ by (2.6). Let
$W=W_{q}(F)\subset F^{*}$
be the subgroup of $q$-Weil numbers and $J\subset W$ be the subgroup generated by $\{j_{d}(\mathbf{a})\mid\mathbf{a}\in A_{d}^{(n)},n\geq 2\}$. Since $\sigma_{h}j_{d}(\mathbf{a})=j_{d}(h\mathbf{a})$, $J$ is $G$-stable as well as $W$. In fact, $J$ is generated by $\{j_{d}(\mathbf{a})\mid\mathbf{a}\in A_{d}^{(2)}\}\cup\{(-1)^{\frac{q-1}{d}}\}$ by (6.3).
@@
which associates to an invertible motive $M$ its Frobenius eigenvalue in $\Lambda\simeq\mathrm{End}(M)$. As a result of Theorem 1.2 and Proposition 8.2, we obtain the following.
@@
Assume that $\mathfrak{Fr}$ is injective. Then $\mathbf{Pic}(\mathbf{Chow}(\kappa,F))_{\mathbb{Q}}$ is isomorphic to $W_{\mathbb{Q}}$ and is generated as a $\mathbb{Q}$-vector space by the classes of Fermat motives $h(F_{d}^{(2)})^{{\boldsymbol{\chi}}}$ (${\boldsymbol{\chi}}\in\mathfrak{X}_{d}^{(2)}$).